OpenStudy (anonymous):

Help! Write answer in standard form. I don't know how to subtract these. (6x-4x^2+4) - (4x^2+5x-2+4x^3)

OpenStudy (anonymous):

Just subtract the variables that have the same bases. eg. 6x-5x and -4x^2-4x^2. You should end up with x-8x^2-4x^3+6. 6x-5x=x -4x^2-4x^2=-8x^2 -4x^3=-4x^3 4-(-2)=6

OpenStudy (mathstudent55):

Remove the parentheses from the trinomial (at left). Then distribute the negative through the entire second parentheses. Every sign inside the second parentheses will change. Then combine like terms.

OpenStudy (anonymous):

so is the first response wrong?

OpenStudy (anonymous):

and @mathstudent55 could you kindly tell me what the final answer would be as I would like to be able to check my solution?